A cube made of lead measures 3 cm on each side and a mass of 310.50 grams. What is it's density in g/cm3?


1
Expert's answer
2021-03-08T08:25:19-0500
"\rho=\dfrac{m}{V}=\dfrac{310.50\ g}{(3\ cm)^3}=11.5\ \dfrac{g}{cm^3}."